# Quantum Models of Consciousness from a Quantum Information Science Perspective

### v1

@@ IV Quantum Consciousness Emerging from the Molecular Interactions Among Neurons
The optimal candidate for a neural qubit in our brain should have a nuclear spin of $\frac{1}{2}$ to ensure long-lived coherence. Therefore, phosphorus emerges as the best candidate for a neural qubit.
@@
producing adenosine monophosphate (AMP) the pyrophosphate ion (PPi). Occasionally, further hydrolysis results in the production of two separate phosphate ions.
@@
When two Posner molecules attempt to bind, their pseudospin states become entangled, regardless of whether the binding is successful. This entanglement is crucial for quantum computations. The two entangled phosphorus nuclear spins, located in different Posner clusters, influence the firing of corresponding neurons during binding reactions. This means that the reactions in two different neurons are entangled.
Furthermore, the Posner cluster has been studied experimentally to gather evidence supporting the potential demonstration of this theory [76, 77, 78, 79, 80]. Theoretical studies have also been conducted to model the system using a Hamiltonian framework and incorporating environmental noise and some quantum information measures such as concurrence. These studies model Posner molecules with six interacting spins, one for each phosphate molecule. The spin Hamiltonian describes these interactions, considering the symmetry of the Posner molecule and examining various configurations. They use concurrence to analyze entanglement within the spin system [81, 82, 83]. Another study approaches it from a quantum computing and information perspective, interpreting each step in Fisher’s proposal as a quantum computing operation, though it does not involve environmental factors [84].
@@ V Study of the entanglement preservation
In our approach, we treat each atom within a phosphate molecule as an identical spin having a magnitude of 1/2. In the context of a phosphate molecule, a cluster of five spins consists of a central spin (associated with the phosphorus atom) surrounded by buffer spins (corresponding to oxygen atoms). The buffer spins interact with individual thermal baths, while the central spin remains isolated from direct environmental interactions. When exposed to a magnetic field along the $z$-axis, an energy difference arises between the lower state (spin-up) and the upper state (spin-down), allowing for a two-level description. The buffer spins eventually reach thermal equilibrium with the environment at an inverse temperature $\beta$. In our study, we analyze the geometry of buffer networks characterized by different arrays of coupling constants, which are either $g$ or 00. These configurations can be represented using planar graphs. Specifically, for a given number of buffer spins, $N$, we explore all feasible buffer networks that can be embedded in a plane. We focus on two extreme cases: one where there is no connectivity among the buffer spins and another where there is maximum connectivity within the buffer spin network.
@@
The Pauli spin-$1/2$ operators for the $i^{th}$ spin are denoted by $\sigma_{x}^{(i)}$, $\sigma_{y}^{(i)}$, and $\sigma_{z}^{(i)}$, and the interaction strength between the spin pair $(i,j)$ is represented by $g_{ij}$. The central spin, labeled by "1", is coupled to the buffer spins at strength $g_{1j}=g\neq 0$ in all geometries under consideration. On the other hand, each buffer network is defined by a different array of coupling constants consisting of zero or $g$ values, which can be represented by a planar graph. In other words, we consider two different geometries of the spins in the buffer network: either they do not interact with each other and are coupled only to the central spin (forming a topology that resembles a star graph), or they are fully connected to all the other spins in the system (giving rise to a complete graph). To be precise, for $N=5$ the fully connected geometry is not a complete graph, as the central spin is positioned right in the middle of the bulk network (more details are given in [28]), hindering the interaction between spin 2 and spin 6, which are uncoupled.
@@
where the unitary contribution to the dynamics is provided by the self-Hamiltonian of the system given in Eq. (10). By assuming weakly coupled buffer spins, local thermal dissipation channels are described by the dissipator in Eq. (11) [86, 87, 88],
@@ VI.1 Hamiltonian Transformation
We now examine the system with fully connected buffer network, and more specifically its Hamiltonian, from a different perspective, in an attempt to provide a theoretical explanation of better entanglement preservation in the fully connected geometry. By applying a unitary transformation to the Hamiltonian, we can transition from a fully connected graph of buffer spins to the other extreme, where the buffer spins do not interact with each other. These fictitious non-interacting buffer spins will be referred to as the dressed buffer spins and the new representation after the transformation as the dressed picture.
We rewrite the Hamiltonian in Eq.(10) in terms of ladder operators as
@@
For $N=2,3,4$ this is equivalent to finding the eigenvalues and eigenvectors of the complete graph $A_{\text{com}}u_{j}=\mu_{j}u_{j}$. We readily obtain $\mu_{1}=\ldots=\mu_{N-1}=-1$ with multiplicity $N-1$, and $\mu_{N}=N-1$ with multiplicity 1. The corresponding eigenvectors are $u_{j}=\frac{1}{\sqrt{2}}\left(\ket{N}-\ket{j}\right)$ for $j=1,\ldots,N-1$, and $u_{N}=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\ket{j}.$ The matrix $D$ in (22) is diagonal and contains the frequencies of the dressed buffer spins, which, following our discussion on the eigenvalues of the complete graph and some trivial algebra for the geometry with $N=5$, are given by:
@@
As discussed before, $\sigma_{1}$ which represents the spin operator of the central spin, remains unchanged.
Next, we rewrite the interaction between the central and the buffer spins in the dressed picture $\tilde{\Sigma}$, making use of the transformation
@@
We can thus define a new vector of interactions $\tilde{v}=(\tilde{g}_{12},\dots,\tilde{g}_{1{N+1}})$. Note, however, that the coupling between the central spin and the buffer spins can be written as $G\sigma_{1}^{+}\sum_{j=2}^{N+1}\sigma_{j}^{-}+H.c.$, and in the case of a complete graph the sum $\sum_{j=2}^{N+1}\sigma_{j}^{-}$ corresponds to a single eigenvector of the graph ($u_{N}$ defined above, with eigenvalue $N-1$). In other words, for $N=2,3,4$ the central spin is coupled to a single dressed buffer spin with energy $\omega+(N-1)G$. Furthermore, the coupling constant is equal to $G\sqrt{N-1}$. Finally, for the case with $N=5$ the buffer spins do not form a perfectly complete graph, as the interaction between spin 2 and spin 6 is missing. However, a similar reasoning applies and it can be shown that the central spin is coupled to only two dressed buffer spins, as depicted in Figure 2 (column C). We remark that the decoupling of the central spin from most of the dressed buffer spins is due to the system Hamiltonian being highly symmetric.
@@ VI.3 Interpretation of the Results
In the non-interacting case, there are $N$ interactions between the central spin and buffer spins (column A in Figure 2). When interactions between buffer spins are introduced (column B in Figure 2), the interaction between the central spin and the dressed buffer spins is reduced to a single interaction for $N=2$ to $N=4$, and to two interactions for $N=5$ (as shown in column C in Figure 2). Using the dressed picture, we can more easily compare these scenarios and understand why interacting buffer spins offer better quantum information protection. In column B, where the buffer spins are maximally interacting, the interaction between the central spin and the buffer network (which is connected to the environment and responsible for the loss of quantum information) is significantly modified. While in the case of non-interacting buffer spins (column A) the central spin is weakly coupled to $N$ independent and locally dissipating buffer spins, if the buffer network is fully connected ($N=2,3,4$) then the central spin interacts with a single dressed buffer spin (column C), which is dissipating into different global baths, in a more and more intense way as a function of $N$ (coupling constant going as $\sqrt{N}$).
@@
Focusing now on the dynamics of the central spin only, the difference between the two scenarios described above lies in a transition from a Markovian to a non-Markovian dynamics. Indeed, if dissipation comes from a collection of identical and weakly coupled systems, the dynamics has a more Markovian behavior corresponding to a roughly flat spectral density of the environment (this is an exact result if we consider $N\rightarrow\infty$ non-interacting buffer spins [85]). On the other hand, a stronger coupling with a single dressed buffer spin implies a more and more non-Markovian behavior, as the dressed spin filters the thermal white noise into Lorentzian noise. Furthermore, it is well-known that a non-Markovian dynamics better preserves quantum information, as non-Markovianity induces a backflow of quantum information from the reservoir to the central spin [90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100]. It is thus reasonable to expect that the fully connected buffer network preserves entanglement during a longer time than the non-interacting network, as we indeed observe.
@@
For $N\geq 5$, it is physically impossible to construct a fully connected network, given the position of the central spin (see e.g. [28]), and for instance for $N=5$ the interaction between spin 2 and spin 6 is missing. So, the dressed buffer network does not reduce to a single interaction. Instead, the central spin is coupled to two non-interacting dressed buffer spins, while three more are left uncoupled. Anyway, we can draw similar conclusions as for the case of perfectly fully connected buffer network, i.e., the dynamics will have a more non-Markovian behavior than in the case of non-interacting buffer network, leading to better entanglement preservation. This behavior, however, will be less pronounced than in the case of a fully connected network.
@@ VI.4 Numerical Verification
To verify numerically this observation we can calculate the trace distance measure of non-Markovianity [101]. Trace distance between two density matrices $\rho_{1}(t)$ and $\rho_{2}(t)$ at time $t$ can be calculated as
@@
Figure 3: Trace norm distance, $D$, between $\rho_{1}$ and $\rho_{2}$ over dimensionless time for different numbers of buffer spins: $N=2$, $N=3$, $N=4$ and $N=5$. The buffer spins interact with each other.
VII Conclusion
The brain remains a profoundly mysterious biological entity, linking us to everything around us and shaping our identity and perception of the world, thus forming the foundation of all scientific fields. Understanding the brain is crucial for addressing brain diseases and potentially developing the next generation of artificial intelligence. Numerous questions arise in the literature: What if our perspective of quantum physics itself needs reviewing? What is the role of the observer in the quantum experiment [102]? Is the physics of the mind and the thought of the observer taken into consideration during any experiment [103]? Could consciousness be related to relativity for instance [104]? Is consciousness dependent on life [105], or can an atom be conscious too [106]?
@@ VII Conclusion
L.G. carried out the simulations and prepared the initial draft. O.P. shaped the conceptual framework of the manuscript and developed the proposals discussed in Sections 2 and 3. M.C. proposed and elaborated the theoretical explanation discussed in Section 6. Ö.E.M. contributed to the conceptualization and provided supervision. All authors contributed to the writing and revision of the manuscript.
Acknowledgements.
This work was supported by the Scientific and Technological Research Council of Türkiye (TÜBITAK) under Project Number 123F150. L. G. and Ö. E. M. thank TÜBITAK for their support. M.C. acknowledges funding from the Research Council of Finland through the Centre of Excellence program grant 336810 and the Finnish Quantum Flagship project 358878 (UH), and from COQUSY project PID2022-140506NB-C21 funded by MCIN/AEI/10.13039/501100011033.

### v2

@@ IV Quantum Consciousness Emerging from the Molecular Interactions Among Neurons
The optimal candidate for a neural qubit in our brain should have a nuclear spin of $\frac{1}{2}$ to ensure long-lived coherence. This is because nuclei with $I>\frac{1}{2}$ experience significant decoherence due to electric field interactions, whereas $I=\frac{1}{2}$ spins are primarily affected by weaker magnetic field perturbations, allowing for much longer coherence times [28]. Therefore, phosphorus emerges as the best candidate for a neural qubit.
@@
producing adenosine monophosphate (AMP) and the pyrophosphate ion (PPi). Occasionally, further hydrolysis results in the production of two separate phosphate ions.
@@
When two Posner molecules attempt to bind, their pseudospin states become entangled, regardless of whether the binding is successful. This entanglement is crucial for quantum computations. The two entangled phosphorus nuclear spins, located in different Posner clusters, influence the firing of corresponding neurons during binding reactions. This means that the reactions in two different neurons are entangled. The schematic showing the two phosphates entangled in two different Posner clusters is provided in Figure 3.
Furthermore, experiments on Posner clusters have been conducted to provide insights into their properties, but their coherence has not yet been studied experimentally [92, 93, 94, 95, 96]. Theoretical studies have also been conducted to model the system using a Hamiltonian framework and incorporating environmental noise and some quantum information measures such as concurrence. These studies model Posner molecules with six interacting spins, one for each phosphate molecule. The spin Hamiltonian describes these interactions, considering the symmetry of the Posner molecule and examining various configurations. They use concurrence to analyze entanglement within the spin system [97, 98, 99]. Another study approaches it from a quantum computing and information perspective, interpreting each step in Fisher’s proposal as a quantum computing operation, though it does not involve environmental factors [100].
@@ V Study of the Entanglement Preservation
In our approach, we treat each atom within a phosphate molecule as an identical spin having a magnitude of 1/2. In the context of a phosphate molecule, a cluster of five spins consists of a central spin (associated with the phosphorus atom) surrounded by buffer spins (corresponding to oxygen atoms). The buffer spins interact with individual thermal baths, while the central spin remains isolated from direct environmental interactions. Figure 1 from reference [29] illustrates this concept. When exposed to a magnetic field along the $z$-axis, an energy difference arises between the lower state (spin-up) and the upper state (spin-down), allowing for a two-level description. The buffer spins eventually reach thermal equilibrium with the environment at an inverse temperature $\beta$. In our study, we analyze the geometry of buffer networks characterized by different arrays of coupling constants, which are either $g$ or 00. These configurations can be represented using planar graphs. Specifically, for a given number of buffer spins, $N$, we explore all feasible buffer networks that can be embedded in a plane (Figure 2 from reference [29] illustrates the different planar graphs). We focus on two extreme cases: one where there is no connectivity among the buffer spins, and another where there is maximum connectivity within the buffer spin network.
@@
The Pauli spin-$1/2$ operators for the $i^{th}$ spin are denoted by $\sigma_{x}^{(i)}$, $\sigma_{y}^{(i)}$, and $\sigma_{z}^{(i)}$, and the interaction strength between the spin pair $(i,j)$ is represented by $g_{ij}$. The central spin, labeled by “1”, is coupled to the buffer spins at strength $g_{1j}=g\neq 0$ in all geometries under consideration. On the other hand, each buffer network is defined by a different array of coupling constants consisting of zero or $g$ values, which can be represented by a planar graph. In other words, we consider two different geometries of the spins in the buffer network: either they do not interact with each other and are coupled only to the central spin (forming a topology that resembles a star graph ), or they are fully connected to all the other spins in the system (giving rise to a complete graph). To be precise, for $N=5$, the fully connected geometry is not a complete graph, as the central spin is positioned right in the middle of the bulk network (more details are given in [29]), hindering the interaction between spin 2 and spin 6, which are uncoupled.
@@
where the unitary contribution to the dynamics is provided by the self-Hamiltonian of the system given in Equation (10). By assuming weakly coupled buffer spins, local thermal dissipation channels are described by the dissipator in Equation (11) [102, 103, 104],
@@
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
@@ VI.1 Hamiltonian Transformation
We now examine the system with a fully connected buffer network, and more specifically its Hamiltonian, from a different perspective, in an attempt to provide a theoretical explanation of better entanglement preservation in the fully connected geometry. By applying a unitary transformation to the Hamiltonian, we can transition from a fully connected graph of buffer spins to the other extreme, where the buffer spins do not interact with each other. These fictitious non-interacting buffer spins will be referred to as the dressed buffer spins and the new representation after the transformation as the dressed picture.
We rewrite the Hamiltonian in Equation (10) in terms of ladder operators as
@@
For $N=2,3,4$, this is equivalent to finding the eigenvalues and eigenvectors of the complete graph $A_{\text{com}}u_{j}=\mu_{j}u_{j}$. We readily obtain $\mu_{1}=\ldots=\mu_{N-1}=-1$ with multiplicity $N-1$, and $\mu_{N}=N-1$ with multiplicity 1. The corresponding eigenvectors are $u_{j}=\frac{1}{\sqrt{2}}\left(\ket{N}-\ket{j}\right)$ for $j=1,\ldots,N-1$, and $u_{N}=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\ket{j}.$ The matrix $D$ in (22) is diagonal and contains the frequencies of the dressed buffer spins, which, following our discussion on the eigenvalues of the complete graph and some trivial algebra for the geometry with $N=5$, are given by:
@@
As discussed before, $\sigma_{1}$, which represents the spin operator of the central spin, remains unchanged.
Next, we rewrite the interaction between the central and the buffer spins in the dressed picture $\tilde{\Sigma}$, making use of the transformation,
@@
We can thus define a new vector of interactions $\tilde{v}=(\tilde{g}_{12},\dots,\tilde{g}_{1{N+1}})$. Note, however, that the coupling between the central spin and the buffer spins can be written as $G\sigma_{1}^{+}\sum_{j=2}^{N+1}\sigma_{j}^{-}+H.c.$, and in the case of a complete graph the sum $\sum_{j=2}^{N+1}\sigma_{j}^{-}$ corresponds to a single eigenvector of the graph ($u_{N}$ defined above, with eigenvalue $N-1$). In other words, for $N=2,3,4$, the central spin is coupled to a single dressed buffer spin with energy $\omega+(N-1)G$. Furthermore, the coupling constant is equal to $G\sqrt{N-1}$. Finally, for the case with $N=5$, the buffer spins do not form a perfectly complete graph, as the interaction between spin 2 and spin 6 is missing. However, a similar reasoning applies and it can be shown that the central spin is coupled to only two dressed buffer spins, as depicted in Figure 5 (column C). We remark that the decoupling of the central spin from most of the dressed buffer spins is due to the system Hamiltonian being highly symmetric.
@@ VI.3 Interpretation of the Results
In the non-interacting case, there are $N$ interactions between the central spin and buffer spins (column A in Figure 5). When interactions between buffer spins are introduced (column B in Figure 5), the interaction between the central spin and the dressed buffer spins is reduced to a single interaction for $N=2$ to $N=4$, and to two interactions for $N=5$ (as shown in column C in Figure 5). Using the dressed picture, we can more easily compare these scenarios and understand why interacting buffer spins offer better quantum information protection. In column B, where the buffer spins are maximally interacting, the interaction between the central spin and the buffer network (which is connected to the environment and responsible for the loss of quantum information) is significantly modified. While in the case of non-interacting buffer spins (column A) the central spin is weakly coupled to $N$ independent and locally dissipating buffer spins, if the buffer network is fully connected ($N=2,3,4$), then the central spin interacts with a single dressed buffer spin (column C), which is dissipating into different global baths in a more and more intense way as a function of $N$ (coupling constant going as $\sqrt{N}$).
@@
Focusing now on the dynamics of the central spin only, the difference between the two scenarios described above lies in a transition from a Markovian to a non-Markovian dynamics. Indeed, if dissipation comes from a collection of identical and weakly coupled systems, the dynamics have a more Markovian behavior corresponding to a roughly flat spectral density of the environment (this is an exact result if we consider $N\rightarrow\infty$ non-interacting buffer spins [101]). On the other hand, a stronger coupling with a single dressed buffer spin implies a more and more non-Markovian behavior, as the dressed spin filters the thermal white noise into Lorentzian noise. Furthermore, it is well-known that non-Markovian dynamics better preserve quantum information, as non-Markovianity induces a backflow of quantum information from the reservoir to the central spin [106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116]. It is thus reasonable to expect that the fully connected buffer network preserves entanglement during a longer time than the non-interacting network, as we indeed observe.
@@
For $N\geq 5$, it is physically impossible to construct a fully connected network, given the position of the central spin (see e.g., [29]), and for instance for $N=5$ the interaction between spin 2 and spin 6 is missing. So, the dressed buffer network does not reduce to a single interaction. Instead, the central spin is coupled to two non-interacting dressed buffer spins, while three more are left uncoupled. We can draw similar conclusions as for the case of perfectly fully connected buffer network, i.e., the dynamics will have a more non-Markovian behavior than in the case of non-interacting buffer network, leading to better entanglement preservation. This behavior, however, will be less pronounced than in the case of a fully connected network.
@@ VI.4 Numerical Verification
To numerically verify this observation, we can calculate the trace distance measure of non-Markovianity [117]. Trace distance between two density matrices $\rho_{1}(t)$ and $\rho_{2}(t)$ at time $t$ can be calculated as
@@
Figure 6: Trace norm distance, $D$, between $\rho_{1}$ and $\rho_{2}$ over dimensionless time for different numbers of buffer spins: $N=2$, $N=3$, $N=4$, and $N=5$. The buffer spins interact with each other.
VII Conclusions
While this overview has aimed to summarize the current state of research on the brain and its functions, it cannot fully encompass the breadth and depth of the field. Our focus has been primarily on the biological and cognitive aspects, and we have not delved into the more speculative intersections between brain research and foundational philosophical or physical theories. These areas, though of growing interest, remain largely in the realm of ongoing exploration and debate. The brain, with its complex and still largely mysterious nature, serves as a central link between our perception of the world and our identity. Its intricate functioning underpins much of scientific inquiry, making the study of the brain vital not only for addressing neurological diseases but also for advancing fields like artificial intelligence and even fundamental physics. Recent discussions in the literature suggest that a deeper understanding of brain function may require revisiting foundational questions in quantum physics and relativity. For instance, the role of the observer in quantum experiments remains an open debate, with potential implications for cognition and perception [119, 118, 122, 121]. Some researchers explore whether consciousness could have a physical basis extending beyond biological systems [119, 118], while others investigate whether principles from relativity might provide insights into the nature of thought and awareness [120]. Though these questions are speculative and far from settled, they underscore the importance of interdisciplinary approaches in advancing our understanding of the brain and its connections to more fundamental domains of science.
@@ VII Conclusions
L.G. carried out the simulations and prepared the initial draft. O.P. shaped the conceptual framework of the manuscript and developed the proposals discussed in Sections II and III. M.C. proposed and elaborated the theoretical explanation discussed in Section VI. Ö.E.M. contributed to the conceptualization and provided supervision..
Acknowledgments
This work was supported by the Scientific and Technological Research Council of Türkiye (TÜBITAK) under Project Number 123F150. L.G. and Ö.E.M. thank TÜBITAK for their support. M.C. acknowledges funding from the Research Council of Finland through the Centre of Excellence program grant 336810 and the Finnish Quantum Flagship project 358878 (UH), and from COQUSY project PID2022-140506NB-C21 funded by MCIN/AEI/10.13039/501100011033. We sincerely thank Professor Travis Craddock from the University of Waterloo for his valuable feedback and thoughtful revisions to Section II of this work.